OpenStudy (anonymous):

how would you solve this problem? I'm getting stuck on how I would put it together ; On Tuesday, a local hamburger shop sold a combined total of hamburgers and cheeseburgers. The number of cheeseburgers sold was two times the number of hamburgers sold. How many hamburgers were sold on Tuesday?

OpenStudy (anonymous):

2x+x=y

OpenStudy (anonymous):

x is the hamburgers and y is your total

OpenStudy (anonymous):

cool?

OpenStudy (anonymous):

right thanks but I'm not 100% sure on how you go from there, like if you divide the "2x" with the 240, you get 12 and end up with the equation x+y=120

OpenStudy (anonymous):

would that be the answer or something?

OpenStudy (anonymous):

what are the numbers?

OpenStudy (anonymous):

check the problem, the total between both hamburgers and cheeseburgers is 240, but the cheeseburgers were sold twice as much then the hamburgers.

OpenStudy (anonymous):

so 2x+x = 3x so 3x=240

OpenStudy (anonymous):

how did you get "3x"?

OpenStudy (anonymous):

oh oh!

OpenStudy (anonymous):

you get it now??

OpenStudy (anonymous):

wait no still not sure

OpenStudy (anonymous):

tell me how you got those numbers to get 3x=240

OpenStudy (anonymous):

so 2x+x=y right??

OpenStudy (anonymous):

and y = 240

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so you combine the x's to get 3x

OpenStudy (anonymous):

oh!! right !

OpenStudy (anonymous):

awesome thanks so much!!

OpenStudy (anonymous):

np